OpenStudy (anonymous):

Help!!!!! Whenever I go to make an equation, the button doesn't work. It literally does NOTHING. Is this a common problem?

OpenStudy (anonymous):

@Teddyiswatshecallsme

OpenStudy (anonymous):

Hi, I tried using it and it is working for me. Maybe you can refresh your browser and try it again. If that still didn't work, I think it's time for you to learn Latex :)

OpenStudy (anonymous):

What is Latex?

OpenStudy (anonymous):

Typesetting program. Say for example, if you'll type `\(\sf \large a^2 +b^2 =c^2\)` the output will be \(\sf \large a^2 + b^2 = c^2\) If you want to learn more about it, you go to Latex Practicing Section :)

OpenStudy (just_one_last_goodbye):

Tbh the Equation tool is just a short cut to LaTeX ^_^ thats why it posts the formula into the comment box.

OpenStudy (usukidoll):

LaTeX is fun to learn...though I have to admit that when I first started using the program, I had to memorize a bunch of commands for the math to display correctly.